# ClimaQA: An Automated Evaluation Framework for Climate Foundation Models

### v1

@@ Abstract
  1. The use of foundation models in climate science has recently gained significant attention. However, a critical issue remains: the lack of a comprehensive evaluation framework capable of assessing the quality and scientific validity of model outputs. To address this issue, we develop ClimaGen (Climate QA Generator), an automated algorithmic framework that generates question-answer pairs from graduate textbooks with climate scientists in the loop. As a result, we present ClimaQA-Gold, an expert-annotated benchmark dataset alongside ClimaQA-Silver, a large-scale, comprehensive synthetic QA dataset for climate science. Finally, we develop evaluation strategies and compare different Large Language Models (LLMs) on our benchmarks. Our results offer novel insights into various approaches used to enhance climate foundation models.
11footnotetext: Department of Computer Science and Engineering, University of California San Diego, La Jolla, California, USA22footnotetext: Halıcıoğlu Data Science Institute, University of California San Diego, La Jolla, California, USA33footnotetext: Scripps Institution of Oceanography, University of California San Diego, La Jolla, California, USA44footnotetext: Department of Linguistics, University of California San Diego, La Jolla, California, USAfootnotetext: Preprint. Under Review. Correspondence to: Taylor Berg-Kirkpatrick <tberg@ucsd.edu>.
@@ 1 Introduction
In recent years, foundation models have attracted significant interest in climate science due to their potential to transform how we approach critical challenges such as climate predictions and understanding the drivers of climate change (Thulke et al., 2024; Nguyen et al., 2024; Cao et al., 2024). However, while these models are powerful, they often fall short when it comes to answering technical questions requiring high precision such as What is the net effect of Arctic stratus clouds on the Arctic climate? Even advanced models like GPT-4 exhibit epistemological inaccuracies in Climate Question-Answering (QA) tasks (Bulian et al., 2024), raising concerns about their reliability in scientific workflows. This highlights the need for a domain-specific evaluation framework to assess the quality and validity of outputs generated by these models. Current benchmarks for Large Language Models (LLMs) predominantly focus on linguistic accuracy or general factual correctness, but they fail to address the unique demands of climate science, where factual rigor, domain-specific knowledge, and robust reasoning are essential. Although some work has explored the scientific evaluation of LLMs, they either rely heavily on manual expert input or employ fully synthetic question generations using LLMs.
@@
To address this issue, we develop ClimaGen, an adaptive learning framework for creating benchmarks in collaboration with domain experts to evaluate scientific foundation models, specifically for climate science but adaptable to other scientific disciplines, shown in Figure 1. This enables us to achieve a balance between utilizing the efficiency of LLMs and the expertise of domain specialists. Using our framework, we introduce a novel benchmark for evaluating foundation models in climate science across three scientific QA task forms: multiple-choice, freeform, and cloze. The questions are designed with varying levels of complexity, challenging the models to demonstrate a range of reasoning abilities from basic factual recall to scientific reasoning and scenario applications. The benchmark consists of two datasets: ClimaQA-Gold – an expert-annotated dataset with a total of 502 questions validated by climate scientists, ensuring high scientific rigor, and ClimaQA-Silver – a large-scale synthetic dataset consisting of 3000 questions generated by our framework, providing substantial ground truth data for model fine-tuning at scale. Together, these datasets enable comprehensive performance assessment of LLMs in climate science, specifically for scientific QA tasks. We evaluate several LLMs on our benchmark under different settings. We observe that most models struggle with reasoning-based multiple-choice questions (MCQs) and Retrieval Augmented Generation (RAG) (Lewis et al., 2020) proves beneficial across different tasks, substantially improving performance as compared to Continued Pre-training and Supervised Fine-tuning.
@@
  • Creation of publicly releasable datasets: expert-annotated (ClimaQA-Gold) and synthetic (ClimaQA-Silver), facilitating both rigorous assessment and large-scale fine-tuning on 3 scientific QA task forms: multiple-choice, freeform, and cloze, with varying levels of complexity.
  • Development of a generalized adaptive learning framework for creating scientific benchmarks at scale in collaboration with domain experts for evaluation of foundation models on scientific accuracy.
  • Evaluation of state-of-the-art foundation models on climate science QA tasks, with insights into improving scientific accuracy.
2 Related Work
@@
ClimaQA-Gold Climate Grad Text 502
@@
ScienceQA (Lu et al., 2022) contains a vast collection of multimodal MCQs manually curated from high school textbooks. Pira2 (Pirozelli et al., 2024) consists of expert-created questions derived from research articles focused on oceans, the Brazilian coast, and climate change. The creation of these benchmarks required substantial manual effort. SciQA (Auer et al., 2023) innovatively generates freeform QA pairs by leveraging hand-crafted queries on the Open Research Knowledge Graph (Jaradeh et al., 2019) primarily drawing from computer science literature. Although these pairs are factually accurate, they do not include an automatic evaluation method for generated responses. Climate Crisis QA (Zhu and Tiwari, 2024) and SciQAG-24D (Wan et al., 2024) explore synthetic data generations using Large Language models. However, their approaches are prone to suffer from hallucinations and lack of scientific validity. To address this, we introduce a gold-standard dataset, rigorously validated by domain experts, alongside a large-scale silver dataset whose generation process was guided by these expert validation labels. Moreover, existing benchmarks generally focus on a single QA format and lack scientifically aligned evaluation metrics. Our benchmark contains questions of three distinct scientific QA task forms at varying levels of complexity, along with evaluation metrics tailored to them. Table 1 presents a comparison of various scientific benchmarks.
@@
Previous work on automated MCQ generation has focused on selecting keywords and generating distractors based on contextual information. (A. Nwafor and E. Onyenwe, 2021)) utilize traditional NLP techniques, such as TF-IDF, for keyword extraction, while (Mehta et al., 2021) leverage BERT (Devlin, 2018) for summarization and keyword extraction, employing WordNet (Miller, 1995) to generate distractors. (Das et al., 2021) applies RAKE (Rose et al., 2010) for keyword extraction and clustering methods for distractor generation. Other approaches, such as utilizing dependency trees (Afzal and Mitkov, 2014), have also been explored for MCQ creation. These methods typically focus on generating MCQs with single-word answers. However, the recent advancements in LLMs have enabled the creation of more complex MCQs with longer, detailed answer choices. (Meißner et al., 2024) demonstrate the automated generation of self-assessment quizzes using LLMs, while (Hang et al., 2024) explore self-refining prompting techniques for improved MCQ generation. Recent studies, including (Olney, 2023) and (Doughty et al., 2024), suggest that LLMs can generate MCQs comparable to those created by humans, though (Grévisse et al., 2024) emphasize the importance of human oversight to ensure the quality and pedagogical relevance of these questions.
3 ClimaQA - Climate Question Answering Benchmark
The ClimaQA benchmark is built on questions generated from graduate-level climate science textbooks, ensuring alignment with the precise terminology and complex theories of the field. These textbooks provide a reliable source for generating both the expert-validated ClimaQA-Gold dataset and the synthetic ClimaQA-Silver dataset. By leveraging textbook content and combining it with expert review, ClimaQA facilitates rigorous evaluation and fine-tuning of LLMs across freeform, multiple-choice, and cloze question-answering tasks in climate science. Our expert-validated dataset, ClimaQA-Gold, ensures that the generated questions are accurate, relevant, and reflect the current understanding of climate science.
3.1 Scientific Question Answering
To thoroughly evaluate a foundation model’s ability to handle scientific questions, we create our benchmark dataset to focus on the different complexities of scientific reasoning. The aim is to test the model’s ability to engage with scientific concepts at different levels of understanding and scenario application.
@@ 3.1 Scientific Question Answering
Together, the benchmark as shown in Table 2 provides a robust framework for evaluating a foundation model’s proficiency in scientific reasoning, critical thinking, and applying knowledge in unseen scenarios.
@@
3.2 Evaluation Metrics
Although assessing multiple-choice question-answering is relatively simple, the other two tasks present more challenges. To address this, we propose and validate the following evaluation metrics for freeform and cloze question-answering. We use these metrics to report experimental results in Section 5
@@ 3.2.1 Freeform QA
Various metrics are employed to evaluate sentence similarity, ranging from surface-level comparisons to deeper semantic analysis. Lexical metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) focus on exact word or n-gram matching, rendering them useful for tasks where token overlap is crucial, such as machine translation and summarization. In contrast, semantic metrics like BERTScore (Zhang et al., 2019), Word Mover’s Distance (WMD) (Huang et al., 2016), and Sentence-BERT (Reimers, 2019) are more advanced, capturing the meanings of sentences through embeddings. These metrics are better suited for tasks that necessitate an understanding of meaning, such as paraphrase detection. However, they may not adequately assess factual accuracy.
@@
To measure the factual accuracy of generated answers relative to reference answers, we propose the use of a factual entailment classifier, reporting the confidence level as the factual accuracy score. Instruction-tuned models, such as GPT-4, have demonstrated superior performance on textual entailment tasks and have shown the ability to generalize across various datasets (Sanyal et al., 2024). We employ the GPT-4o-mini model with the prompt above for factual entailment. This method achieved $81\%$ zero-shot classification accuracy on the Climate-Fever dataset (Diggelmann et al., 2020) indicating the ability to measure factual accuracy.
To use this for scoring freeform QA, the reference answer was used as the evidence and the generated answer was used as the claim. The confidence score was computed by applying sigmoid smoothing to the logit scores, with a temperature parameter set to $T=5$. Note that the choice of $T$ does not alter the score trend; it was selected to optimally scale values between 0 and 1.
Factual $\text{Factual Accuracy}=\text{SigmoidSmooth}\left(\frac{p_{s}}{p_{s}+p_{r}}\right)$
@@ 3.2.2 Cloze QA
To evaluate the robustness of this approach, we synthetically generated 219 cloze questions with our framework (described in Section 4), which were answered by the GPT-4o-mini model. We collected 32 questions where the generated answers did not exactly match the reference answer. These answers were then labeled as wrong or correct by domain experts based on scientific and contextual accuracy. We plotted the average cosine similarities of phrases for these questions, concluding that a context window of size 4 most effectively differentiates between correct and incorrect answers. The cosine similarities were subsequently rescaled to emphasize these differences.
@@ 4 ClimaGen - Automated Benchmark Creation
In this section, we describe ClimaGen as shown in Figure 1, the framework used to create the ClimaQA dataset from Climate Textbooks. By leveraging RAG and prompt engineering, we systematically generate and refine questions of varying complexity levels. Domain expert annotators ensure quality, producing a semi-synthetic benchmark for the evaluation of AI systems in complex scientific inquiry. Additionally, we automate annotation by fine-tuning an LLM on human-labeled data, enabling the creation of a large-scale synthetic dataset for fine-tuning tasks. The techniques discussed here could be generalized to aid in the semi-automatic production of benchmarks for other scientific fields.
@@ 4.1 Textbook Dataset
LLMs are typically pre-trained on extensive general internet data, which often contains noise and misinformation. This limitation is particularly significant in fields like climate science, where a precise understanding of specialized terminology and concepts is crucial. To evaluate LLMs’ proficiency in climate science, we employed graduate-level climate science textbooks as a reliable source of specialized knowledge, providing accurate scientific information that better represents the technical terms and nuanced theories integral to this discipline. We collected 18 textbooks and extracted their content, which was preprocessed to ensure cleanliness and relevance, making it suitable for downstream applications such as benchmark creation, continuous pre-training, and RAG. Specifically, a held-out representative set of 5 textbooks was used for the benchmark creation process.
@@ 4.3 Domain Expert Annotation
To mitigate this, we developed an interactive web application that enables climate scientists to review and annotate the generated questions. For freeform and MCQs, scientists validated the correctness of the content, while for cloze questions, they selected which scientific term to mask, ensuring alignment with scientific standards. By combining human expertise with AI, we curated 181 freeform, 161 MCQ, and 160 cloze questions, forming the ClimaQA-Gold dataset, reviewed and validated by domain experts.
@@ 4.4 Automated Annotation
We observed that around 85% of multiple-choice questions (MCQs) and around 90% of freeform question-answers (QAs) were valid, indicating high-quality question generation. Experiments across different train-test splits show that the evaluator models enhance the quality of the generated MCQ question set by 10% and the Freeform question set by 5% as shown in Appendix A.4.1. Additionally, we fine-tune a separate LLM to mark scientific terms from given statements as shown in Appendix A.4.2, automating the cloze annotation process. Using this framework, we generated 1000 freeform QAs, 1000 MCQs, and 1000 cloze questions, collectively forming the ClimaQA-Silver dataset, produced synthetically at scale without manual intervention.

### v2

# ClimaQA: An Automated Evaluation Framework for Climate Question Answering Models

@@ Abstract
The use of Large Language Models (LLMs) in climate science has recently gained significant attention. However, a critical issue remains: the lack of a comprehensive evaluation framework capable of assessing the quality and scientific validity of model outputs. To address this issue, we develop ClimaGen (Climate QA Generator), an adaptive learning framework that generates question-answer pairs from graduate textbooks with climate scientists in the loop. As a result, we present ClimaQA-Gold, an expert-annotated benchmark dataset alongside ClimaQA-Silver, a large-scale, comprehensive synthetic QA dataset for climate science. Finally, we develop evaluation strategies and compare different LLMs on our benchmarks. Our results offer novel insights into various approaches used to enhance knowledge of climate LLMs. ClimaQA’s source code is publicly available at https://github.com/Rose-STL-Lab/genie-climaqa
@@ 1 Introduction
Climate change is one of the most pressing global challenges today, with profound impacts on ecosystems, economies, and societies. In recent years, Large Language Models (LLMs) have gained significant interest in climate science (Thulke et al., 2024; Nguyen et al., 2024; Cao et al., 2024) due to their potential to transform climate predictions and enable applications in climate policy analysis, environmental decision-making, and public education. By improving LLMs’ understanding of climate science, we can empower stakeholders to make informed decisions, develop actionable solutions, and foster broader awareness of climate issues. However, while LLMs are powerful, they often fall short when it comes to answering technical questions requiring high precision such as What is the net effect of Arctic stratus clouds on the Arctic climate? Even advanced models like GPT-4 exhibit epistemological inaccuracies in Climate Question-Answering (QA) tasks (Bulian et al., 2024), raising concerns about their reliability in scientific workflows.
This highlights the need for a domain-specific evaluation framework to assess the quality and validity of outputs generated by these models. Current benchmarks for Large Language Models (LLMs) predominantly focus on linguistic accuracy or general factual correctness (Bai & Wang, 2021), but they fail to address the unique demands of climate science, where factual rigor, domain-specific knowledge, and robust reasoning are essential. Although some work has explored the scientific evaluation of LLMs (Table 1), they either rely heavily on manual expert input or employ fully synthetic question generations.
@@
To address this issue, we develop ClimaGen, an adaptive learning framework for creating benchmarks in collaboration with domain experts to evaluate scientific question-answering models, specifically for climate science but adaptable to other scientific disciplines, shown in Figure 1. This enables us to achieve a balance between utilizing the efficiency of LLMs and the expertise of domain specialists.
Using our framework, we introduce a novel benchmark for evaluating question-answering models in climate science across three scientific QA task forms: multiple-choice, freeform, and cloze. The questions are designed with varying levels of complexity, challenging the models to demonstrate a range of reasoning abilities from basic factual recall to scientific reasoning and scenario applications. The benchmark consists of two datasets: ClimaQA-Gold – an expert-annotated dataset with a total of 566 questions validated by climate scientists, ensuring high scientific rigor, and ClimaQA-Silver – a large-scale synthetic dataset consisting of 3000 questions generated by our framework, providing substantial ground truth data for model fine-tuning at scale. Together, these datasets enable comprehensive performance assessment of LLMs in climate science, specifically for scientific QA tasks. We evaluate several LLMs on our benchmark under different settings. We observe that most models struggle with reasoning-based multiple-choice questions (MCQs) and Retrieval Augmented Generation (RAG) (Lewis et al., 2020) significantly outperforms Continued Pre-training and Supervised Fine-tuning across different tasks.
@@
  • Creation of publicly releasable datasets: expert-annotated (ClimaQA-Gold) and synthetic (ClimaQA-Silver), along with tailored evaluation metrics facilitating both rigorous assessment and large-scale fine-tuning on 3 scientific QA task forms: multiple-choice, freeform, and cloze, with varying levels of complexity.
  • Development of a generalized adaptive learning framework (ClimaGen) for creating scientific benchmarks at scale in collaboration with domain experts for evaluation of natural language question-answering models on scientific accuracy.
  • Evaluation of state-of-the-art LLMs on climate science QA tasks, with insights into improving scientific accuracy.
@@
ClimaQA-Gold Climate Grad Text 566
@@
2 Related Work
ScienceQA (Lu et al., 2022) contains a vast collection of multimodal MCQs manually curated from high school textbooks. Pira2 (Pirozelli et al., 2024) consists of expert-created questions derived from research articles focused on oceans, the Brazilian coast, and climate change. The creation of these benchmarks required substantial manual effort. SciQA (Auer et al., 2023) innovatively generates freeform QA pairs by leveraging hand-crafted queries on the Open Research Knowledge Graph (Jaradeh et al., 2019) primarily drawing from computer science literature. Although these pairs are factually accurate, they do not include an automatic evaluation method for generated responses. Climate Crisis QA (Zhu & Tiwari, 2024) and SciQAG-24D (Wan et al., 2024) explore synthetic data generations using Large Language models. However, their approaches are prone to suffer from hallucinations and lack of scientific validity. To address this, we introduce a gold-standard dataset, rigorously validated by domain experts, alongside a large-scale silver dataset whose generation process was guided by these expert validation labels. Moreover, existing benchmarks generally focus on a single QA format and lack scientifically aligned evaluation metrics. Our benchmark contains questions of three distinct scientific QA task forms at varying levels of complexity, along with evaluation metrics tailored to them. Table 1 presents a comparison of various scientific benchmarks.
Previous work on automated MCQ generation has focused on selecting keywords and generating distractors based on contextual information. (A. Nwafor & E. Onyenwe, 2021)) utilize traditional NLP techniques, such as TF-IDF, for keyword extraction, while (Mehta et al., 2021) leverage BERT (Devlin, 2018) for summarization and keyword extraction, employing WordNet (Miller, 1995) to generate distractors. (Das et al., 2021) applies RAKE (Rose et al., 2010) for keyword extraction and clustering methods for distractor generation. Other approaches, such as utilizing dependency trees (Afzal & Mitkov, 2014), have also been explored for MCQ creation. These methods typically focus on generating MCQs with single-word answers. However, the recent advancements in LLMs have enabled the creation of more complex MCQs with longer, detailed answer choices. (Meißner et al., 2024) demonstrate the automated generation of self-assessment quizzes using LLMs, while (Hang et al., 2024) explore self-refining prompting techniques for improved MCQ generation. Recent studies, including (Olney, 2023) and (Doughty et al., 2024), suggest that LLMs can generate MCQs comparable to those created by humans, though (Grévisse et al., 2024) emphasize the importance of human oversight to ensure the quality and pedagogical relevance of these questions.
3 ClimaQA - Climate Question Answering Benchmark
The ClimaQA benchmark is built on questions generated from graduate-level climate science textbooks, ensuring alignment with the precise terminology and complex theories of the field. These textbooks provide a reliable source for generating both the expert-validated ClimaQA-Gold dataset and the synthetic ClimaQA-Silver dataset. By leveraging textbook content and combining it with expert review, ClimaQA facilitates rigorous evaluation and fine-tuning of LLMs across freeform, multiple-choice, and cloze question-answering tasks in climate science. Our expert-validated dataset, ClimaQA-Gold, ensures that the evaluation questions are accurate, relevant, and reflect the current understanding of climate science.
3.1 Scientific Question Answering
To thoroughly evaluate a model’s ability to handle scientific questions, we create our benchmark dataset to focus on the different complexities of scientific reasoning. The aim is to test the model’s ability to engage with scientific concepts at different levels of understanding and scenario application.
@@ 3.1 Scientific Question Answering
Together, the benchmark as shown in Table 2 provides a robust framework for evaluating an LLM’s proficiency in scientific reasoning, critical thinking, and applying knowledge in unseen scenarios.
3.2 Evaluation Metrics
Although assessing multiple-choice question-answering is relatively simple, the other two tasks present more challenges. To address this, we propose and validate the following evaluation metrics for freeform and cloze question-answering. A more detailed case study to demonstrate the robustness of these metrics can be found in the Appendix A.2.3. We use these metrics to report experimental results in Section 5
@@ 3.2.1 Freeform QA
Various metrics are employed to evaluate sentence similarity, ranging from surface-level comparisons to deeper semantic analysis. Lexical metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) focus on exact word or n-gram matching, rendering them useful for tasks where token overlap is crucial, such as machine translation. In contrast, semantic metrics like BERTScore (Zhang et al., 2019), Word Mover’s Distance (WMD) (Huang et al., 2016), and Sentence-BERT (Reimers, 2019) are more advanced, capturing the meanings of sentences through embeddings. These metrics are better suited for tasks that necessitate an understanding of meaning, such as paraphrase detection. However, they may not adequately assess factual accuracy.
To measure the factual accuracy of generated answers relative to reference answers, we propose the use of a factual entailment classifier, reporting the confidence level as the factual accuracy score. Instruction-tuned models, such as GPT-4, have demonstrated superior performance on textual entailment tasks and have shown the ability to generalize across various datasets (Sanyal et al., 2024). We employ the GPT-4o-mini model with the prompt below for factual entailment. This method achieved $81\%$ zero-shot classification accuracy on the Climate-Fever dataset (Diggelmann et al., 2020) indicating the ability to measure factual accuracy.
@@
To use this for scoring freeform QA, the reference answer was used as the evidence and the generated answer was used as the claim. The confidence score was computed by applying sigmoid smoothing to the logit scores, with the temperature parameter set to $T=5$. Note that the choice of $T$ does not alter the score trend; it was selected to optimally scale values between 0 and 1. If $l_{s}$ and $l_{r}$ are the logit scores for SUPPORTS and REFUTES respectively, then
Factual $\text{Factual Accuracy}=\text{SigmoidSmooth}\left(\frac{l_{s}}{l_{s}+l_{r}}\right)$
@@ 3.2.2 Cloze QA
To evaluate the robustness of this approach, we synthetically generated 219 cloze questions with our framework (described in Section 4), which were answered by the GPT-4o-mini model. We collected 32 questions where the generated answers did not exactly match the reference answer. These answers were then labeled as wrong or correct by domain experts based on scientific and contextual accuracy. We plotted the average cosine similarities of phrases for these questions as shown in Figure 4, concluding that a context window of size 4 most effectively differentiates between correct and incorrect answers. This configuration yields the maximum difference in scores while maintaining sufficiently high scores for correct answers. The cosine similarities were subsequently rescaled to emphasize these differences. If $e_{1}$ and $e_{2}$ are the embeddings of the respective phrases as mentioned above, then
@@ 4 ClimaGen - Automated Benchmark Creation
In this section, we describe ClimaGen as shown in Figure 1, the framework used to create the ClimaQA dataset from Climate Textbooks. By leveraging RAG and prompt engineering, we systematically generate and refine questions of varying complexity levels. Domain expert annotators ensure quality, producing a semi-synthetic benchmark for the evaluation of AI systems in complex scientific inquiry. Additionally, we automate annotation by fine-tuning an LLM on human-labeled data, enabling the creation of a large-scale synthetic dataset for fine-tuning tasks. The techniques discussed here could be generalized to aid in the semi-automatic production of benchmarks for other scientific fields aswell.
@@ 4.1 Textbook Dataset
LLMs are typically pre-trained on extensive general internet data, which often contains noise and misinformation. This limitation is particularly significant in fields like climate science, where a precise understanding of specialized terminology and concepts is crucial. To evaluate LLMs’ proficiency in climate science, we employed graduate-level climate science textbooks as a reliable source of specialized knowledge, providing accurate scientific information that better represents the technical terms and nuanced theories integral to this discipline. We collected 18 textbooks (Table 4) that broadly represent a mixture of graduate and expert literature on the physical climate with a particular focus on the role of aerosol in the climate system - one of the critical sources of uncertainty in climate projections. The content was extracted and preprocessed to ensure cleanliness and relevance, making it suitable for downstream applications such as benchmark creation, continuous pre-training, and RAG. A held-out set of 5 textbooks (Figure 6), carefully selected to represent varying levels of technical and qualitative depth across a broad range of key topics in climate science, was utilized for the benchmark creation process.
@@ 4.3 Domain Expert Annotation
To mitigate this, we developed an interactive web application that enables climate scientists to review and annotate the generated questions. For freeform and MCQs, the scientists validated the correctness of the content, while for cloze questions, they selected which scientific term to mask, ensuring alignment with scientific standards. They also identified common reasons for rejecting generated QA pairs during validation, providing valuable insights into improvement of the QA generation framework as discussed in Appendix A.5.1. By combining human expertise with AI, we curated 245 freeform, 161 MCQ, and 160 cloze questions, forming the ClimaQA-Gold dataset, reviewed and validated by domain experts.
@@ 4.4 Automated Annotation
We observed that around 85% of multiple-choice questions (MCQs) and around 90% of freeform question-answers (QAs) were valid, indicating high-quality question generation. Experiments across different train-test splits show that the evaluator models enhance the quality of the generated MCQ question set by 10% and the Freeform question set by 5% as shown in Appendix A.5.2. Additionally, we fine-tune a separate model to mark scientific terms from given statements as shown in Appendix A.5.3, automating the cloze annotation process. Using this framework, we generated 1000 freeform QAs, 1000 MCQs, and 1000 cloze questions, collectively forming the ClimaQA-Silver dataset, produced synthetically at scale without manual intervention.